OpenStudy (anonymous):

Evaluate x/y for x = 2/3 and y = 6/7 I know its substituting, but come someone explain ? :3

OpenStudy (anonymous):

can*

OpenStudy (amorfide):

You want x/y if you have x=2/3 y=6/7 do x divided by y

OpenStudy (amorfide):

divide the numbers

OpenStudy (nincompoop):

yeah so you place the values of x and then you do the same for y (2/3)/(6/7) if you took division of fraction in 4th grade, you would know that dividing two fractions is basically obtaining the reciprocal of one of the fractions then multiply it to the other fraction so it becomes

OpenStudy (nincompoop):

2/3 * 7/6

OpenStudy (anonymous):

K thanks :)